# Fast LISA likelihood approximations by downsampling

### v1

@@ Abstract
The LISA gravitational wave observatory is due to launch in 2035. Researchers are currently working out the data analysis picture, including determining requirements of, and prospects for, parameterised gravitational wave models used in Bayesian inference, of which early stage inspirals of low and intermediate mass black hole binaries constitute a significant and important part. With datasets consisting of up to hundreds of millions of datapoints, their likelihood functions can be extremely expensive to compute. We present an approximation procedure for accurately reproducing the likelihood function (from time-domain data) for such signals in simulated environments, significantly reducing this cost. The method is simply to discard most of the datapoints and define a new inner product operator in terms of the remaining data, which closely resembles the original inner product on the model parameter space. Highly accurate reproductions of the likelihood can be achieved using just a few hundred to a few thousand datapoints, which can reduce parameter estimation (PE) time by factors of up to tens of thousands (compared to estimated frequency domain PE time), however, the end user must be responsible for ensuring convergence. The time-domain implementation is particularly useful for modelling waveform modifications arising from non-trivial astrophysical environments. The contents of this article provide the theoretical basis of the software package dolfen.
@@ 1 Introduction
In order to analyse very large, purely simulated datasets, given limited computing power and time, we investigate a means of reducing the computational costs without jeopardising the accuracy of the results. Our approach will be to discard the majority of the datapoints and modify properties of the remaining data, or the methods used to analyse or transform it, in order to compensate for the original information content of interest. We shall refer to this procedure as downsampling, and in particular, we aim to ascertain the robustness, limitations and benefits of downsampling a given dataset, such that the likelihood function used in Bayesian parameter estimation (PE) will be minimally affected. With this view, the task is an optimisation problem. Using fewer data points generally means fewer numerical operations are required to evaluate the likelihood function at a given point in the parameter space. We will see that downsampling allows a drastic reduction in the likelihood computation time, by factors ranging from hundreds to hundreds of thousands in the extreme cases. The particular focus here is the study of ‘slowly evolving’ gravitational wave (GW) signals that we expect to obtain from the Laser Interferometer Space Antenna (LISA) detector [1] (with special emphasis on compact binary inspirals), but the methods employed here are general and may be used with other datasets of signals that can be said to be slowly evolving in time. To our knowledge, the downsampling techniques we introduce here for PE have not previously been considered.
@@
The inspiral stage of black hole binaries (BHBs) may last millions of years before merger, where the vast majority of the duration of GW emission constitutes a very slowly evolving signal of gradually increasing amplitude and frequency (a characteristic ‘chirp’ [2]). The LISA detector will be sensitive to a good portion of the lower frequency ($\sim$1–100 mHz) part, and may be active for $\sim$5–10 years [1]. For the most slowly evolving signals hovering near the upper frequency sensitivity of $\sim$100 mHz throughout LISA’s lifetime, we may realistically expect datasets consisting of $\sim 10^{9}$ data points at the Nyquist rate.
Such large datasets require a great deal of computing resources for PE, but in preparatory stages of research when working with entirely simulated datasets, we are free to use any sort of technique to acquire the information we desire, including redefining the signal and/or detector model. Our sole objective is the rapid procurement of an accurate representation of the likelihood function, and downsampling makes use of the liberties of simulated environments to achieve this very cheaply. We show that downsampling can accurately reproduce the results one can expect to obtain from a real-world analysis of a real-world dataset, in, relatively, very short time. For real-world scenarios, this procedure is not applicable since we have no control of either the signal or noise models; this procedure is only applicable in simulations and will therefore serve only to provide experimental insight into which sorts of analyses of large data sets could be fruitful, the results we can expect to see when real-world analyses are performed on the real-world dataset, and signal modelling requirements, for example.
It is instructive to place this method and its results in the context of related works that take other approaches in either approximating, or in some other way speeding up the evaluation of PE analyses. Perhaps the most frequently employed formalism is the Fisher information approximation; see for example [3, 4, 5, 6, 7, 8]. This formalism provides a multivariate Gaussian approximation of the likelihood that gives a lower bound in the standard error on the parameters given by the proper posterior. This can be of some use, but in many cases is a poor approximation of the posterior; for a detailed discussion of why this is not generally reliable (except in some particular regimes), see [9, 10] for further details. Other methods that exploit common properties of the waveforms under study, by finding ways to cheaply produce waveform templates, include reduced order quadrature (ROQ) [11], heterodyning/relative binning [12], and so-called adaptive frequency resolution [13]. Another related technique: Template-Interpolation [14], entails computing the frequency domain waveform excluding in those parts which are known to be roughly linear, and interpolating over those gaps in the waveform between the points that are explicitly computed.
Downsampling is not similar to any of these techniques. We shall of course describe downsampling in detail; the definitions of signal/noise/inner product operations, and whatever else that is required, are modified such that the remaining datapoints constrain the signal parameters just as well as the fully sampled dataset would have (within some reasonable threshold of variation of the morphology of the posteriors). A useful way to think of downsampling is from the point of view of information theory. A posterior has a given morphology or ‘shape’ (considering isoprobability contours, for example). If an optimal encoding scheme was found requiring $E$ nats (natural information units) to describe this morphology, then $E$ is called the entropy of the posterior. We neither know a posterior nor its optimal encoding scheme a priori, when given data and signal/noise models. However, the signal model is an encoding scheme (lossless); the signal model using fewer datapoints is a lossy, but presumably a well-informed encoding scheme. Our downsampling procedure attempts to tweak this lossy encoding scheme to minimise the divergence of the representation of the posterior.
@@
It is also appropriate to comment briefly on the choice of deriving this approximation for time-, rather than frequency-domain signals. A great deal of signal analysis, particularly in the field of GW data analysis, is performed in the frequency-domain; the inner product of frequency-domain vectors is cheaper to compute since datapoints are often uncorrelated there. For a significant subset of the signals expected to be detected by LISA, however, parameterised models (or modifications of, from, for example, time-delays in the GW induced by galactic orbital motion of binaries and environmental effects) are straightforward to write down in the time-domain, and conversely very difficult to formulate in the frequency-domain. With this in mind, and the fact that some real LISA analyses are likely to take place in the time-domain, it will be useful to have a time-domain method. This will also be useful for developing understanding of the intricacies of LISA time-series signal modelling and data analysis.
@@
This paper is structured as follows. In Section 2, we introduce the framework in which to define the problem and other useful results that will be required subsequently. We then define and develop different approaches to downsampling in Section 3, discussing some of the concepts and problems that must be considered for a faithful reproduction of the LISA posterior. In Section 4, details of how high-dimensional posteriors can be numerically compared and downsampling acceptance criteria are derived. We then detail the physical set up of the problem and compute and compare posteriors to evaluate downsampling procedures experimentally. Section 5 presents the projected improvements in the rate of evaluation of PE attained by downsampling, and gives closing remarks. The contents of this article constitute the theoretical basis of the software package dolfen.
@@ 2 Preliminaries
2.1 Slow evolution waveforms
Downsampling cannot be expected to work for analysis of any given parameterised waveform. An important requirement for analysis of a waveform to be amenable to downsampling is that it must be slowly evolving. We loosely define a waveform here as being slowly evolving over a time interval $\mathcal{S}$ if the average value of each Fisher Information Matrix (see Section 3.2) element, at time $t$ over ‘intermediate duration’ time intervals $\mathcal{I}_{t}$ (with a range of order $\sim$100 times the reciprocal of the Nyquist rate of the waveform in $\mathcal{S}$) centred at $t$ is approximately constant for all $t\in\mathcal{S}$. Alternatively, for our GW waveforms, we could define slow evolution as having roughly constant coalescence phase marginalised Fisher matrix elements of each time-series datapoint of the signal. A definition of this sort is required in order to be able to classify sinusoidal waveforms as being slowly evolving if their frequency and amplitude parameters are roughly constant, despite their oscillating nature; their Fisher information matrix elements also oscillate as functions of time, and taking the average value over intermediate time intervals (or marginalising over phase) ‘smooths out’ the oscillations.
To see why slow evolution is an important requirement for downsampling, consider the simplest time-series signal that is clearly slowly evolving: a constant signal. One could take datapoints from any location; these would provide roughly equal information on the constant. Conversely, at the other extreme, consider a signal with an instantaneous step from one constant to another at a certain time; one should clearly classify this as not slowly evolving. The datapoints just before and after the step provide the most information on the location of the step. Downsampling in this case will fail; removing any datapoint eliminates the ability to determine the original likelihood of the step location between the remaining samples, at the same resolution. Since we aim to introduce a general, signal independent downsampling procedure, we therefore restrict applications to signals of slow time evolution.
@@ 2.2 Bayesian inference
If a signal contains a waveform generated by some parameterised model, we can estimate those generating parameters using methods of Bayesian inference as is now commonplace in GW data analysis. We define the posterior probability distribution of model parameters using Bayes’ theorem,
$p(\bm{\theta}\,|\,\textbf{d})=\frac{p(\bm{\theta})\,p(\textbf{d}\,|\,\bm{\theta})}{p(\mathbf{d})}\,,$ (1)
@@
We may define our own prior based on expectations derived from other means; for the present study we will mostly choose uniform priors with a suitably chosen window centred on the true parameters. The remaining factor to be determined is the likelihood function. Assuming a Gaussian noise model, we can write the probability of some noise realisation $\mathbf{n}$, a vector of densities, occurring (i.e., the likelihood) as
@@
[…]
@@ 3.1.1 Sample selection schemes
  • Hybrid sampling - combining uniform and random sampling; roughly half samples taken uniformly, and half taken randomly from original dataset.
One sample selection scheme may allow recovery of a fully sampled posterior in fewer samples than some other sampling scheme. Attempts to optimise the downsampling procedure using specific schemes should likely need to be studied for each type of model (even for each dataset considered); the efforts required for this will almost certainly outweigh any benefits and shall not be considered here. However, since we must choose some scheme, we will perform a comparison of various schemes using some fiducial system to decide, empirically, on the best performing scheme.
The scheme of greatest interest is random sampling. It is expected to give a good representation of the signal for high enough sampling density. Uniform (/periodic/regular) downsampling is expected to lead to aliasing and related effects. By randomly sampling, this problem can be mostly eliminated [16], so long as the number of samples used is not catastrophically low, causing the structure of the posterior to break down. In some preliminary tests, however, it was noticed that uniform downsampling is very effective: the structure of the posterior remains stable down to a surprisingly low sampling rate. We surmise that this might be due to the fact that, as a result of the non-white noise profile, the strain value of any given sample is informed by a number of its neighbouring samples, which are located at a distance so as to at least capture the Nyquist frequency components of the signal. Thus despite the vast downsampling of uncorrelated samples, some high frequency information remains since the decorrelation process encodes it in the residual.
Because uniform sampling performs well, we will examine the hybrid scheme, however, uniform sampling alone shall not be considered due to its potential to compromise the integrity of the likelihood in a general setting, by the reasons outlined above. The hope is that incorporating some degree of uniformity in the datapoint selection will ensure evenness in the relative spread of information across the signal and be less prone to datapoint selection mishaps, securely allowing fewer samples to retain the required information, whilst effects such as aliasing are suppressed by the random samples. For these reasons we shall also examine the performance of cluster sampling.

### v2

@@ Abstract
The Laser Interferometer Space Antenna (LISA) is due to launch in the mid-2030s. A key challenge for LISA data analysis is efficient Bayesian inference with parametrised gravitational-wave models, particularly for early inspirals of low- and intermediate-mass black-hole binaries, where time series can contain $\sim 10^{8}$$10^{9}$ samples and naive likelihood evaluations become prohibitively expensive. For purely simulated studies, we present a simple time-domain likelihood-approximation scheme for such signals. The method retains only a small subset of samples and defines a modified noise-weighted inner product on this reduced set that closely reproduces the original inner product on the waveform manifold. Because this alters the effective noise model, the scheme is intended for simulated LISA data rather than direct analysis of real LISA data. In our examples, the resulting posteriors agree to high accuracy with those obtained using much denser sampling, while retaining only $10^{3}$$10^{4}$ samples. The computational cost then scales linearly with the number of retained samples $N_{\mathrm{s}}$, so the speed-up is roughly $N/N_{\mathrm{s}}$, where $N$ is the original data length; for realistic LISA-like datasets with $N\sim 10^{8}$$10^{9}$ this would correspond to gains of order $10^{4}$$10^{6}$ over straightforward frequency-domain likelihood evaluations of the full dataset. The time-domain formulation is particularly convenient for incorporating waveform modifications from non-trivial astrophysical environments or alternative-gravity effects. The results presented here provide the theoretical basis for the software package Dolfen.
@@ 1 Introduction
Motivated by the need to perform Bayesian parameter estimation (PE) on very large, purely simulated datasets with limited computing resources, we investigate a means of reducing the computational costs without jeopardising the accuracy of the results. Our approach will be to discard the majority of the data points and redefine a noise-weighted inner product on the retained data points (typically retaining only $N_{\mathrm{s}}\sim 10^{3}$$10^{4}$ samples) to compensate for the discarded information. We shall refer to this procedure in this paper as downsampling, and aim to quantify the robustness, limitations and benefits of downsampling using a given dataset, in particular such that the likelihood function used in Bayesian PE is minimally affected. From this perspective, the task is an optimisation problem. Using fewer data points generally means fewer numerical operations are required to evaluate the likelihood function at a given point in parameter space. We will see that downsampling allows a drastic reduction in the likelihood computation time. The particular focus here is the study of ‘slowly evolving’ gravitational wave (GW) signals that we expect to obtain from the Laser Interferometer Space Antenna (LISA) detector [LISAwhitepaper] (with special emphasis on compact binary inspirals), but the methods employed here are general and may be used with other datasets of signals that can be said to be slowly evolving in time.
@@
The inspiral stage of black hole binaries (BHBs) may last millions of years before merger, where the vast majority of the duration of GW emission constitutes a very slowly evolving signal of gradually increasing amplitude and frequency (a characteristic ‘chirp’ [chirp]). The LISA detector will be sensitive to a good portion of the lower frequency ($\sim$1–100 mHz) part, and may be active for $\sim$5–10 years [LISAwhitepaper]. For the most slowly evolving signals hovering near the upper frequency sensitivity of $\sim$100 mHz throughout LISA’s lifetime, we may realistically expect datasets consisting of $\sim 10^{8}$$10^{9}$ samples at the Nyquist rate.
Such large datasets require substantial computing resources for PE, but in preparatory work with fully simulated datasets we are free to use any approach that reproduces the desired information, including redefining aspects of the signal and/or detector model. Our sole objective is to obtain an accurate representation of the likelihood function as cheaply as possible, which downsampling achieves very effectively. This scheme is not intended for direct use with real LISA data, because it alters the effective noise model in a way that is only justified for simulated data with known noise properties; its use is therefore limited to fully simulated PE environments. We use downsampling here as a fast yet high accuracy exploratory tool: to assess the viability of proposed analysis strategies, to quantify the waveform-modelling accuracy required for LISA, and to study the features of large-dataset likelihoods.
It is instructive to place this method and its results in the context of related PE acceleration approaches. Perhaps the most frequently employed formalism is the Fisher information approximation; see for example [GrimmHarms, MBinGWTC1, FundPhyswGWs, ExpndLISAfrmGrnd, LISAEMRIsFisher, TransOrbRes]. This formalism provides a multivariate Gaussian approximation of the likelihood, giving a lower bound on the standard error of the parameters under the proper posterior. In many cases it is a poor approximation of the posterior; for a detailed discussion of why this is not generally reliable (except in some particular regimes), see [UseAbuseFisher, InadequateFIM]. In follow-up work, we will show that this problem becomes severe for LISA likelihood functions of extended GW waveform models.
Other methods that exploit common properties of the waveforms under study, by finding ways to cheaply produce waveform templates, include reduced order quadrature (ROQ) [ROQ], heterodyning/relative binning [cornish2013fastfishermatriceslazy, RelativeBinning], and so-called adaptive frequency resolution (AFR) [AdaptFreqRes]. The related technique of Template-Interpolation [TemplateInterpolation] entails computing the frequency-domain waveform excluding those parts which are known to be roughly linear, and interpolating over the gaps between explicitly computed points. While AFR and template interpolation operate in the frequency domain, ROQ relies on detector- and prior-specific offline training to construct reduced bases, and heterodyning makes very efficient use of a high-overlap reference waveform to compress the likelihood evaluation. Each of these approaches has regimes where it is highly effective, but they typically require problem-specific preparation (basis construction, training, or reference selection) when the waveform family, priors, or data model are changed.
Even in the presence of highly optimised heterodyned pipelines, a time-domain downsampling scheme remains useful. Heterodyning and relative binning achieve large gains by expanding around a high-overlap reference waveform, which makes them extremely efficient once a suitable reference has been chosen but also ties the compressed data product to that specific waveform family and signal model. By contrast, the downsampling developed here is defined entirely at the level of the data and the likelihood inner product: one selects a subset of samples with associated weights and then evaluates arbitrary time-domain waveforms only on that subset. The same reduced data and inner product can be reused for different waveform models and alternative-gravity signals, without regenerating model-dependent heterodyned streams each time the signal model changes. The accuracy of the approximation is controlled by a single parameter $N_{\mathrm{s}}$, which can be increased until the resulting posterior stabilises, as quantified in Section 4.2. In this sense, downsampling does not compete with heterodyning as a niche optimisation, but provides a generic, reusable data-compression layer that is valuable even when heterodyned likelihoods are available; there may even be scope to combine the two approaches by applying downsampling to heterodyned data products, though we shall not pursue this here.
A useful corollary is that nothing in the construction assumes a fixed number of sources. Once a sample subset and weight vector have been chosen for a given data stream, any composite waveform built from that stream is treated in exactly the same way: whatever configuration the sampler proposes is evaluated, whitened, and inserted into the same inner product on the retained samples. The computational bookkeeping associated with adding or removing sources is therefore decoupled from the data-compression step. We do not explore such applications in this work, but they provide a natural direction for future studies.
A demanding and increasingly routine task in GW astronomy is to quantify bias from waveform incompleteness [capuano2025, Kapil2024]. Signals are injected that include extra physics (e.g. eccentric periastron bursts, self-force terms, non-GR dispersion) and recovery is attempted with a baseline model that omits those effects. Each injection requires a full Bayesian run, and many are needed to chart bias versus SNR and source parameters. In such studies it is advantageous to use accelerators that do not require retraining or recalibration when the injected physics or recovery model are changed. This makes downsampling a natural candidate tool for large-scale bias surveys. The computational cost scales with the number of retained samples, and, where desired, the accuracy of the approximation can be checked by repeating the analysis for two or more values of $N_{\mathrm{s}}$ and confirming that the resulting posteriors are consistent within sampling noise. The implementation is minimal, compatible with gapped data, and integrates cleanly with standard samplers.
@@
It is also appropriate to comment briefly on the choice of deriving this approximation for time-, rather than frequency-domain signals. A great deal of signal analysis, particularly in the field of GW data analysis, is performed in the frequency domain; for stationary Gaussian noise the covariance is diagonal there, and FFT-based likelihood evaluations are correspondingly cheap. For a significant subset of the signals expected to be detected by LISA, however, parameterised models (or modifications thereof, from, for example, time-delays in the GW induced by galactic orbital motion of binaries and environmental effects) are straightforward to write down in the time domain, and conversely very difficult to formulate in the frequency domain. With this in mind, and the fact that some real LISA analyses are likely to take place in the time domain, it is very useful to have a time-domain method. This will also be useful for developing understanding of the intricacies of LISA time-series signal modelling and data analysis. It has recently been pointed out to the author that downsampling has been used in the recent work of Ref. [speri]. In that paper, the authors applied downsampling on strongly Gaussian posteriors of high-SNR EMRI waveforms in the frequency domain. In this paper, downsampling is somewhat more generalised, applying also to low-SNR signals with highly structured likelihood functions, and the time-domain formulation can often be very useful for constructing intricate and comprehensive waveform models.
In this work we formulate a generic time-domain downsampling scheme based on a modified noise-weighted inner product on a retained subset of samples; demonstrate on LISA-like inspiral signals that full posteriors can be recovered to high accuracy using only $N_{\mathrm{s}}\sim 10^{3}$$10^{4}$ samples, with speed-ups of order $10^{4}$$10^{6}$ over straightforward frequency-domain likelihood evaluations of the full dataset; introduce posterior-distance diagnostics to quantify the discrepancy between downsampled posteriors and to define practical convergence criteria; and provide an implementation in the software package Dolfen that fits directly into standard data analysis pipelines, enabling these techniques to be used in broader simulation studies.
@@
This paper is structured as follows. In Section 2 we introduce the framework used to define the problem and collect several results that will be required later. In Section 3 we define and develop different approaches to downsampling, discussing some of the concepts and issues that must be addressed to obtain a faithful reproduction of the LISA posterior. Section 4 introduces numerical diagnostics for comparing high-dimensional posteriors and derives practical acceptance criteria for downsampling schemes; in that section we also describe test cases and compute and compare posteriors to evaluate the procedures experimentally. Section 5 presents the projected improvements in the rate of evaluation of PE attained by downsampling and summarises our main conclusions about the performance and limitations of the method. Finally, Section 6 describes the Dolfen software package, which implements the techniques developed here and is intended for use in broader simulation studies.
@@ 2 Preliminaries
2.1 Slowly evolving waveforms
Downsampling cannot be expected to work for an arbitrary parameterised waveform. An important requirement for a waveform to be amenable to downsampling is that it must be ‘slowly evolving’. We loosely define a waveform as slowly evolving over a time interval $\mathcal{S}$ if, for each element of the Fisher information matrix (see Section 3.2), the average value over a moving window $\mathcal{I}_{t}$ centred at time $t$ is approximately constant for all $t\in\mathcal{S}$. Here $\mathcal{I}_{t}$ is an ‘intermediate-duration’ interval whose length is of order $10^{2}$ times the sampling interval (i.e. the reciprocal Nyquist rate) of the waveform on $\mathcal{S}$. For the GW waveforms considered here, an equivalent characterisation is that the Fisher matrix elements, marginalised over coalescence phase, vary only slowly as a function of time sample across $\mathcal{S}$. A definition of this sort is needed to classify sinusoidal waveforms as slowly evolving when their frequency and amplitude parameters are roughly constant: although the waveform oscillates and the pointwise Fisher matrix elements oscillate with it, averaging over intermediate windows (or marginalising over phase) smooths out these oscillations.
To see why slow evolution is an important requirement for downsampling, consider the simplest time-series signal that is clearly slowly evolving: a constant signal. Taking data points from any location would provide equal information on the constant. Conversely, consider a signal with an instantaneous step from one constant to another at a certain time; one should clearly classify this as not slowly evolving. The data points just before and after the step provide the most information on the location of the step. Downsampling in this case will fail; removing any data point eliminates the ability to determine the original likelihood of the step location between the remaining samples, at the same resolution. Since we aim to introduce a general, signal-independent downsampling procedure, we restrict applications to signals of slow time evolution.
@@ 2.2 Bayesian inference
If the data contain a signal described by a parametrised waveform model, the associated parameters can be inferred using Bayesian methods, as is now standard in GW data analysis. The posterior probability density for the parameters $\bm{\theta}$ given the data $\mathbf{d}$ is defined by Bayes’ theorem,
$p(\bm{\theta}\,|\,\mathbf{d})=\frac{p(\bm{\theta})\,p(\mathbf{d}\,|\,\bm{\theta})}{p(\mathbf{d})}\,$, (1)
@@
For simplicity, we take uniform priors with a suitably chosen window centred on the true parameters, and appropriate sinusoidal priors for angular parameters. The remaining factor to be determined is the likelihood function. Assuming a Gaussian noise model, we can write the probability of some noise realisation $\mathbf{n}$ occurring (i.e., the likelihood) as
@@
where $\mathbf{C}$ is the (noise) covariance matrix, $N\equiv\dim(\mathbf{n})$, and ($\langle\cdot|\cdot\rangle$) is an inner product. In the time-domain this can be written [Cornish_2020]
@@ 3.1.1 Sample selection schemes
  • Hybrid sampling - combining uniform and random sampling; roughly half the samples taken uniformly, and half taken randomly from original dataset.
One sample selection scheme may allow recovery of a fully sampled posterior in fewer samples than some other scheme. Attempts to optimise the downsampling procedure using specific schemes should likely need to be studied for each type of model (even for each dataset considered); the efforts required for this will almost certainly outweigh any benefits and shall not be considered here. However, since we must choose some scheme, we will perform a comparison of various schemes using some fiducial system to decide, empirically, on the best performing scheme.
The scheme of greatest interest is random sampling, which is expected to give a good representation of the signal for high enough sampling density. Uniform (/periodic/regular) downsampling is expected to lead to aliasing and related effects, but by randomly sampling, this problem can be mostly eliminated [Bretthorst], so long as the number of samples used is not catastrophically low, causing the structure of the posterior to break down. In some preliminary tests, however, it was noticed that uniform downsampling is very effective: the structure of the posterior remains stable down to a surprisingly low sampling rate. We surmise that this might be due to the fact that, as a result of the non-white noise profile, the strain value of any given sample is informed by a number of its neighbouring samples, which are located at a distance so as to at least capture the Nyquist frequency components of the signal. Thus despite the vast downsampling of uncorrelated samples, some high frequency information remains since the decorrelation process encodes it in the residual.
Because uniform sampling performs well, we will examine the hybrid scheme, however, uniform sampling alone shall not be considered due to its potential to compromise the integrity of the likelihood in a general setting, by the reasons outlined above. The idea is that incorporating some degree of uniformity in the data point selection will ensure evenness in the relative spread of information across the signal and be less prone to random data point selection problems, securely allowing fewer samples to retain the required information, whilst effects such as aliasing are suppressed by the random samples. For these reasons we shall also examine the performance of cluster sampling.